OpenStudy (4815162342):

Assuming all volume measurements are made at the same temperature and pressure, how many liters of water vapor can be produced when 8.23 liters of oxygen gas react with excess hydrogen gas? Show all of the work used to solve this problem. Balanced equation: 2H2 (g) + O2 (g) yields 2H2 O(g)

OpenStudy (4815162342):

@Matt6288

OpenStudy (matt6288):

2H2 (g) + O2 (g) ---> 2H2O (g) 1 mol of O2 yields 2 mol H2O. Likewise, 1 L of O2 yields 2 L H2O. So, 8.23 L O2 yields 16.46 L H2O.

OpenStudy (4815162342):

What would be the equation?

OpenStudy (4815162342):

Are you still typing the reply?

OpenStudy (4815162342):

Can you help me find the right answer?

OpenStudy (matt6288):

Volume of gases are equal if there moles are equal 1 mole of O2 makes 2 moles of water vapor So volume of water vapor would be twice the volume of O2 Volume of water vapor=8.23* 2 Volume of water vapor=16.46 liters 8.23LO2 x 2 molesH2O/1 mole O2

OpenStudy (4815162342):

Is the answer 16.46? @Matt6288

OpenStudy (4815162342):

@sweetburger

OpenStudy (matt6288):

yes

OpenStudy (4815162342):

How should i show my work??

OpenStudy (4815162342):

@sweetburger is this right?

OpenStudy (sweetburger):

Looks good to me.

OpenStudy (matt6288):

This is how you can show your work: 8.23LO2 x 2 molesH2O/1 mole O2=16.46L of water vapor